Brian and his sister, Linda, agreed that they would each mow one-half of the lawn. The lawn is 25 feet by 45 feet rectangle. the lawn mower cuts a path two feet wide. If Brian starts at one corner and mows a path completely around the outside, how many times should he go around to mow one half of the lawn? Be sure to include your calculations for the entire yard and each trip Brian will take. I'm not even sure where to start!
Answer by stanbon(75887)

Brian and his sister, Linda, agreed that they would each mow one-half of the lawn.
Start by drawing a picture of the lawn.
------------
The lawn is 25 feet by 45 feet rectangle. the lawn mower cuts a path two feet wide.
Label the picture with this information.
------------
If Brian starts at one corner and mows a path completely around the outside, how many times should he go around to mow one half of the lawn? Be sure to include your calculations for the entire yard and each trip Brian will take.
=====================================================
Total area is 45*25 = 1125 sq ft. ; half of that area is 562.5 sq. ft
----------------------------------
1st time around area: (45+21+45+21)*2 = 264 sq ft.
---
The remaining uncut area is a 41 by 21 rectangle
---
2nd time around area: (41 + 17 + 41 + 17)*2 = 232 sq. ft.
---
The remaining uncut area is a 37 by 14 rectangle
Brian's remaining area to cut is 562.5-(264+232) = 66.5 sq. ft.
------
By now you should see how to finish the problem.
Use your picture and keep track of the areas.
==================================================
Cheers,
Stan H.